# Compressing Scene Dynamics: A Generative Approach

### v1

@@ Abstract
This paper proposes to learn generative priors from the motion patterns instead of video contents for generative video compression. The priors are derived from small motion dynamics in common scenes such as swinging trees in the wind and floating boat on the sea. Utilizing such compact motion priors, a novel generative scene dynamics compression framework is built to realize ultra-low bit-rate communication and high-quality reconstruction for diverse scene contents. At the encoder side, motion priors are characterized into compact representations in a dense-to-sparse manner. At the decoder side, the decoded motion priors serve as the trajectory hints for scene dynamics reconstruction via a diffusion-based flow-driven generator. The experimental results illustrate that the proposed method can achieve superior rate-distortion performance and outperform the state-of-the-art conventional video codec Versatile Video Coding (VVC) on scene dynamics sequences. The project page can be found at https://github.com/xyzysz/GNVDC.
1 Introduction
@@
Motion is the eternal melody of the world. Efficiently characterizing motion patterns and reducing temporal redundancy is of vital importance to video coding techniques. With the recent advancements of “Artificial Intelligence Generated Content (AIGC)” techniques, generative video coding (GVC) is proposed to reduce the transmission redundancy and achieve ultra-low bit-rate coding by leveraging the statistical regularities in particular video contents such as human faces [1]. Specifically, the motion flow can be predicted from compact feature representations that are extracted from the key frame and inter frames. Then, the strong generation ability of the deep generative model can guarantee visual-pleasing reconstruction by animating the key frame with the corresponding optical flows [2]. Such a design can warrant superior Rate-Distortion (RD) performance and outperform Versatile Video Coding (VVC) with a large margin in terms of perceptual metrics [3].
@@
Existing GVC methods usually use explicit representations with physical meaning as the precursors of motion representations. For example, Deep Animation Codec [4] adopts 2D key-points representations [5] to build generative video-conferencing codec, Face Video-to-Video Synthesis [6] further leverages 3D key-points for free-view talking-head synthesis, and an extreme generative human-oriented video codec [7] is built by compressing the articulated motion representations [8]. Meanwhile, implicit features are also explored for direct motion representations. Compact Feature Temporal Evolution [9] encapsulates temporal trajectories into $4\times 4$ matrices, and Latent Image Animation [10] extracts weighting coefficients for decomposed motion vectors. However, these methods exploit prior knowledge from video contents under the same scenario, hindering their trained models to generalize to more diverse video contents.
@@
Recently, Generative Image Dynamics [11] illustrates the effectiveness of learning the distribution of natural motions conditioned on a given image. It can generate natural oscillatory motions across scenes using the Diffusion Model [12]. Connecting this motion prior with the prior-based generative video coding raises an intriguing question: Can we characterize specific motion pattern priors into suitable representations that are independent of video contents, so as to enable generative video coding to generalize to a wider range of scenes? To verify this idea, in this paper, we follow [11] and choose the most common motion patterns, i.e., small motion dynamics in everyday scenes, to explore the possibility of building Generative Scene Dynamics Coding framework (Dynamics-Codec). Herein, we limit “small motion dynamics” to the minor, rigid and non-articulated motion in ubiquitous daily scenes, such as oscillations and rectilinear movements of objects. It is expected that this Dynamics-Codec framework can directly extract compact feature representations from motion flows instead of video contents and realize generative reconstruction for various kinds of dynamic scenes. The difference between the video-content-prior-based and motion-pattern-prior-based generative video coding is illustrated in Figure 1.
(a) Video-content-prior-based GVC.
(b) Motion-pattern-prior-based GVC.
@@
Figure 1: Comparison of the video-content-prior-based and motion-pattern-prior-based generative video coding.
@@
The efforts to establish the proposed Dynamics-Codec are two-fold. First, compact motion tokens are identified from motion flows with a motion tokenizer in a dense-to-sparse manner for efficient compression. To benefit the modelling of certain motion patterns, the main characteristics of moving areas are first identified by motion sampling from the dense motions to form sparse motions. Then, the spare motions are further condensed into compact tokens. Second, a robust decoder is designed to generate reconstructions with key frame and motion information across different video contents. Herein, the Stable Video Diffusion (SVD) model [13] that is pre-trained on a wide range of data distribution is utilized to enable the generalizability to various scenes, and a pre-trained Control-Net like motion flow adaptor [14] is adopted to ensure the seamless flow-driving control in a feature-warping manner with the recovered dense motions.
The contribution of this paper can be summarized as the following,
@@
  • The experiment results show that the proposed Dynamics-Codec can achieve superior rate-distortion performance as well as subjective quality against the state-of-the-art conventional video codec VVC.
2 The Proposed Dynamics-Codec
@@
2.1  The Dynamics-Codec Framework
Figure 2: The detailed structure of proposed Dynamics-Codec. The networks in orange blocks are optimized from scratch, and the networks in blue block is pre-trained.
The detailed structure of proposed Dynamics-Codec is shown in Figure 2. At the encoder side, the key frame (i.e., the first frame of the video sequences) is compressed by conventional VVC codec and transmitted as image bit-stream. For inter frames, dense optical flows are first extracted between the key frame and every subsequent inter frame. Then, each dense flow is sampled by a watershed-based method [15, 16] to obtain the sparse motion and the corresponding motion mask, which are further down-sampled and vectorized to form motion tokens. To further eliminate the coding redundancy, all tokens are inter-predicted with the adjacent token, and the quantized residuals are encoded by Context Adaptive Binary Arithmetic Coding (CABAC).
At the decoder side, the reconstructed key frame is decoded from the image bit-stream, and fed into a feature extractor. Meanwhile, the motion tokens are obtained by context-based entropy decoding and feature compensation from the feature bit-stream. The dense motions are subsequently reconstructed by leveraging the internal relationship between the extracted key frame features and the decoded motion tokens. Finally, the recovered dense motions are fed into a flow-driven diffusion-based generator for denoising generation of reconstructed inter frames.
@@ 2.1  The Dynamics-Codec Framework
2.2  Optical Flow Tokenizer Feature representation is essential to generative video coding. Previous GVC works mainly focus on utilizing video content priors for characterizing temporal trajectory on homogeneous data such as human face or human body [2]. In this paper, we advance the idea of video-content prior to motion-pattern prior that requires direct feature extraction from optical flows. To preserve the major motion information and benefit the efficient representation, the motion tokenizer is designed as a dense-to-sparse scheme.
Specifically, dense motion is first extracted with an off-the-shelf and latest motion estimation model Unimatch [17], which utilizes a transformer-based multi-task framework to solve the unified dense correspondence matching problem. Without loss of generality, we denote the key frame and inter frame as I and P with the dimension of $3\times H\times W$, the motion extraction can be denoted as,
$\textbf{f}_{d}=\Theta(\textbf{I},\textbf{P})$ (1)
To select the primary motion trajectories under the consideration of both its appearance and intensity, a topological-distance watershed map sampling [16] with motion edge detection and Non-maximum Suppression [18] is leveraged to obtain the sparse motions and the motion masks, where the latter serves to distinguish between sampled positions [15] with zero-motion and unsampled positions,
$\textbf{f}_{s},\textbf{m}_{s}=\Phi(\textbf{f}_{d})$ (2)
where $\Phi$ denotes motion sampling. $\textbf{f}_{s}\in\mathbb{R}^{H\times W\times 2}$ and $\textbf{m}_{s}\in\mathbb{R}^{H\times W\times 1}$ denote the sparse motion and the motion mask respectively. Then, a sparse motion compressor is utilized to transform sparse motions to compact motion tokens, where the motion feature is first obtained by a flow feature extractor $\textbf{e}_{f}$,
$\textbf{y}_{f}=\textbf{e}_{f}(concat[\textbf{f}_{s},\textbf{m}_{s}])$ (3)
where $\textbf{y}_{f}$ denotes the flow feature and $concat$ denotes the concatenation operation. Herein, this flow feature extractor is designed as a U-Net [19] like structure with up-sampling, down-sampling and short-cut connections. Then, two motion token vectors are derived by two vectorizers $\textbf{v}_{w}$ and $\textbf{v}_{b}$,
$\textbf{w}=\textbf{v}_{w}(\textbf{y}_{f})$ (4)
$\textbf{b}=\textbf{v}_{b}(\textbf{y}_{f})$ (5)
where $\textbf{w}\in\mathbb{R}^{N_{v}\times 1}$ and $\textbf{b}\in\mathbb{R}^{N_{v}\times 1}$. $N_{v}$ denotes the number of tokens in each vector. Here, the vectorizers are designed as cascaded convolutional layers and Generalized Divisive Normalization layers [20]. In this way, high-fidelity dense motions are transformed to very compact motion tokens, which served as high-level semantic motion hints, through the guidance of intermediate sparse motions. Such a process could exploit the compact motion pattern prior in motion flows, benefiting both high compressibility of transmitted features and decent accuracy of subsequent motion recovery.

### v4

# Tokenizing Motion: A Generative Approach for Scene Dynamics Compression

@@ Abstract.
This paper proposes a novel generative video compression framework that leverages motion pattern priors, derived from subtle dynamics in common scenes (e.g., swaying flowers or a boat drifting on water), rather than relying on video content priors (e.g., talking faces or human bodies). These compact motion priors enable a new approach to ultra-low bit-rate communication while achieving high-quality reconstruction across diverse scene contents. At the encoder side, motion priors can be streamlined into compact representations via a dense-to-sparse transformation. At the decoder side, these priors facilitate the reconstruction of scene dynamics using an advanced flow-driven diffusion model. Experimental results illustrate that the proposed method can achieve superior rate-distortion performance and outperform the state-of-the-art conventional video codec Enhanced Compression Model (ECM) on scene dynamics sequences. The project page can be found at https://github.com/xyzysz/GNVDC.
Generative Video Coding, Motion Tokenization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Image compressionccs: Theory of computation Data compression
1. Introduction
@@
Motion is the eternal melody of the world. Efficiently characterizing motion patterns and reducing temporal redundancy is of vital importance to video coding techniques. With the recent advancements of “Artificial Intelligence Generated Content (AIGC)” techniques, generative video coding (GVC) is proposed to reduce the transmission redundancy and achieve ultra-low bit-rate coding by leveraging the statistical regularities in particular video contents such as human faces (gfvc-review). Specifically, the motion flow can be predicted from compact feature representations that are extracted from the key frame and inter frames. Then, the strong generation ability of the deep generative model can guarantee visual-pleasing reconstruction by animating the key frame with the corresponding optical flows (generativevisualcompressionreview). Such a design can warrant superior Rate-Distortion (RD) performance and outperform the latest conventional video coding standard Versatile Video Coding (VVC) with a large margin in terms of perceptual metrics (cttr).
@@
Existing GVC methods usually use explicit representations with physical meaning as the precursors of motion representations. For example, Deep Animation Codec (dac) adopts 2D key-points representations (fomm) to build generative video-conferencing codec, Face Video-to-Video Synthesis (fv2v) further leverages 3D key-points for free-view talking-head synthesis, and an extreme generative human-oriented video codec (mraa-codec) is built by compressing the articulated motion representations (mraa). Meanwhile, implicit features are also explored for direct motion representations. Compact Feature Temporal Evolution (cfte) encapsulates temporal trajectories into $4\times 4$ matrices, and Latent Image Animation (lia) extracts weighting coefficients for decomposed motion vectors. However, these methods exploit prior knowledge from video contents under the same scenario, hindering their trained models to generalize to more diverse video contents.
@@
Recently, Generative Image Dynamics (generative-image-dynamics) illustrate the effectiveness of learning the distribution of natural motions conditioned on a given image. It can generate natural oscillatory motions across scenes using the Diffusion Model (diffusion). Connecting this motion prior with the prior-based generative video coding raises an intriguing question: Can we characterize specific motion pattern priors into suitable representations that are independent of video contents, so as to enable generative video coding to generalize to a wider range of scenes? To verify this idea, in this paper, we follow (generative-image-dynamics; dynamics-dcc) and choose the most common motion patterns, i.e., small motion dynamics in everyday scenes, to explore the possibility of building Generative Scene Dynamics Coding framework (Dynamics-Codec). Herein, we limit “small motion dynamics” to the minor, rigid and non-articulated motion in ubiquitous daily scenes, such as oscillations and rectilinear movements of objects. It is expected that this Dynamics-Codec framework can directly extract compact feature representations from motion flows instead of video contents and realize generative reconstruction for various kinds of dynamic scenes. The difference between the video-content-prior-based and motion-pattern-prior-based generative video coding is illustrated in Figure 1.
(a) Video-content-prior-based.
(b) Motion-pattern-prior-based.
@@
Figure 1. Comparison of the video-content-prior-based and motion-pattern-prior-based generative video coding.
@@
The Dynamics-Codec is developed through two primary efforts. First, a motion tokenizer extracts compact motion tokens from motion flows in a dense-to-sparse manner for efficient compression. To enhance modeling of specific motion patterns, key movement regions are identified through motion sampling, converting dense motions into sparse representations, which are subsequently compressed into compact tokens. Second, a robust decoder reconstructs videos using key frame and motion data across diverse content. Leveraging the pre-trained Stable Video Diffusion (SVD) model (svd), trained on a broad data distribution, ensures generalizability across varied scenes. Additionally, a pre-trained Control-Net-like motion flow adaptor (mofa) enables seamless flow-driven control via feature warping with reconstructed dense motions.
This paper is an extension version of our previous conference paper (dynamics-dcc), which only contains one page length of preliminary illustration of the motion-pattern-prior-based video coding framework. In this paper, we provide a comprehensive discussion on motivation, related works, detailed technical designs, optimization strategies, and extensive experimental validations. The contribution of this paper can be summarized as the following,
@@
  • The experiment results show that the proposed Dynamics-Codec can achieve superior rate-distortion performance as well as subjective quality against the state-of-the-art conventional video codec ECM.
2. Related Work
2.1. Hybrid Video Compression
For over forty years, video coding technologies have undergone remarkable advancements, leading to the development of standardized hybrid codecs with outstanding compression performance, such as Advanced Video Coding (AVC) (avc), High Efficiency Video Coding (HEVC) (hevc), and Versatile Video Coding (VVC) (vvc). The Joint Video Experts Team (JVET), a collaboration between ISO/IEC SC 29 and ITU-T SG16, is currently spearheading efforts to create a next-generation codec that outperforms VVC. This initiative centers on iteratively enhancing coding tools within the Enhanced Compression Model (ECM) reference software (ecm). In parallel, investigations into Neural Network-based Video Coding (NNVC) (nnvc) and the refinement of traditional coding techniques (deep-inloop; deep-intra) aim to push the boundaries of compression efficiency. In this paper, we employ the hybrid codec to encode key frames in video sequences, achieving superior compression efficiency while ensuring high-quality texture references for generating subsequent inter frames.
2.2. End-to-End Video Compression
Unlike hybrid video coding, where coding tools are independently designed and optimized, end-to-end coding models are trained holistically in a data-driven manner. Ballé et al. pioneered this approach in image domain with a transform-quantization-coding pipeline using convolutional neural networks and variational autoencoders (joint; variational; factorized). Inspired by such philosophy, DVC (dvc) marked a breakthrough in end-to-end video coding by implementing all components with deep neural networks. Building on DVC (dvc), DCVC (dcvc) introduced conditional coding in the feature domain, while DCVC-TCM (dcvc-tcm) enhanced compression through temporal context mining. DCVC-HEM (dcvc-hem) incorporated an efficient spatial-temporal entropy model, and DCVC-DC (dcvc-dc) increased context diversity across temporal and spatial dimensions. Recently, DCVC-FM (dcvc-fm) improved quality range and stabilized long prediction chains via feature modulation, outperforming the Enhanced Compression Model (ECM) (ecm) in the Low-Delay-Bidirectional (LDB) configuration. DCVC-RT (dcvc-rt) focused on real-time applications, achieving competitive performance with reduced complexity, providing a more practical solution for neural-network-based video coding.
2.3. Generative Video Compression with Priors
To further boost the rate-distortion performances and go beyond the conventional transform-quantization-coding paradigm, generative video compression (GVC) is proposed to leverage the compact priors of the specific domain and strong generation ability of deep generative models for ultra-low bit-rate coding and perceptually high-quality reconstructions. Specifically, key-frames are compressed by conventional codec while subsequent inter-frames are represented by semantic features. At the decoder side, the key-frames are animated by the motion fields which are derived from the semantic features, to reconstruct inter-frames in a generative manner. Early attempts of GVC mainly focus on evolving the deep image animation model (fomm) into Generative Face Video Coding (GFVC) (gfvc-survey-trans) models by leveraging rich semantic priors of human faces. To achieve that, diverse feature representations are explored including 2D key-points (dac), 3D key-points (fv2v), compact matrices (cfte; cttr). In parallel, hybrid schemes are also incorporated for better reconstruction quality, such as multi-layer coding (hdac), predictive coding (rdac), multi-frame reference (mr-dac), multi-view fusion (multiview), bi-directional prediction (bi-direction), progressive coding (progressive-gfvc), scalable coding (pleno) and interactive coding (ifvc). To develop a more practical GFVC system and collaborate with standardized conventional codecs, JVET has investigated the integration of GFVC techniques into conventional codec with Supplimemtary Enhancement Information (SEI) messages (gfvc-sei-trans).
Subsequently, the GVC framework is extended to more complicated human body contents with articulated movements (mraa). Principle-Component-Analysis-based key-points representation is first explored (mraa-codec), then MTTF (mttf) utilizes multi-granularity features to achieve both precise feature transmission and precision motion recovery, IHVC (ihvc) incorporates interactive semantics for controllable body parts manipulation, and IMT (imt) leverages implicit motion representation with attention-based motion transfer instead of explicit motion field with warpping-based deformation, whose effectiveness has previously been verified on other contents like flowers and foliage (imf). To further expand the dimension of GVC, Sparse2Dense (sparse2dense) facilitates 3D vertices’ prediction from 3D key-points along with human video reconstruction. However, existing GVC methods predominantly focus on video-content-prior-based schemes, which limits their generalizability to other domain’s contents, once trained on a specific domain like human faces or human bodies. Furthermore, these existing GVC schemes mainly utilize Generative Adversarial Networks (GAN) (gan) for reconstruction, whose generation capability has been suppressed by Diffusion Models (diffusion) or Auto-Regressive Models (llm_beates_diffusion) in recent year. In this paper, we attempt to explore motion-pattern-prior-based GVC framework that can generalize to various scene dynamics contents, and leverage the powerful diffusion-based generator for high-quality reconstruction.
2.4. Motion-Driven Image-to-Video Generation
Temporal modeling is an essential for in video generation, and many works attempt to incorporate motion field as an explicit driving condition. Early attempts include animating fluid elements using dense motion predicted from source image and optional arrow-like sparse motion (controllale_fluid; animate_fluid). More recently, the development of Video Diffusion Models (lvdm) has significantly advanced the image-to-video generation field with longer sequence length, higher resolution, better temporal coherence and more diverse controlling signals (diffusion-review). As for motion-driven scenarios, DrugNUWA (dragnuwa) samples sparse trajectories from dense motion and integrate them with source image and text hints into diffusion generator with multi-scale fusion, Motion-I2V (motion-i2v) utilizes two-stage pipeline of text-guided motion generation and motion-guided video generation, and MotionCtrl (motionctrl) decomposes camera motion and objective motion and offers independent control. To further improve the accuracy of temporal alignment, MOFA (mofa) designs a motion adaptor to warp multi-scale image features with dense motion, Tora (tora) introduces temporal attention mechanism to integrate motion patches into Diffusion Transformer (DiT), while DragAnything (draganything) achieves objects-level control with an entity semantic representation. In this paper, we propose to leverage the powerful generation ability of pre-trained motion-driven diffusion model for scene dynamics reconstruction in GVC framework, which makes it possible to explore the prior modelling from motion patterns instead of video contents to enhance the generalizability of GVC between various scenes.
3. Proposed Method
Figure 2. The overall framework of proposed Dynamics-Codec. The networks in orange blocks are optimized from scratch, and the networks in blue block is pre-trained.
3.1. The Dynamics-Codec Framework
The detailed structure of proposed Dynamics-Codec is shown in Figure 2. At the encoder side, the key frame (i.e., the first frame of the video sequences) is compressed by conventional codec and transmitted as image bit-stream. For inter frames, dense optical flows are first extracted between the key frame and every subsequent inter frame. Then, each dense flow is sampled by a watershed-based method (cmp; watershed) to obtain the sparse motion and the corresponding motion mask, which are further down-sampled and vectorized to form motion tokens. To further eliminate the coding redundancy, all tokens are inter-predicted with the adjacent token, and the quantized residuals are encoded by Context Adaptive Binary Arithmetic Coding (CABAC).
At the decoder side, the reconstructed key frame is decoded by conventional codec from the image bit-stream, and fed into a feature extractor. Meanwhile, the motion tokens are obtained by context-based entropy decoding and feature compensation from the feature bit-stream. The dense motions are subsequently reconstructed by leveraging the internal relationship between the extracted key frame features and the decoded motion tokens. Finally, the recovered dense motions are fed into a flow-driven diffusion-based generator for denoising generation of reconstructed inter frames.
3.2. Optical Flow Tokenizer
Figure 3. The detailed structure of motion tokenizer.
Feature representation is essential to generative video coding. Previous GVC works mainly focus on utilizing video content priors for characterizing temporal trajectory on homogeneous data such as human face or human body (generativevisualcompressionreview). In this paper, we advance the idea of video-content prior to motion-pattern prior that requires direct feature extraction from optical flows. To preserve the major motion information and benefit the efficient representation, the motion tokenizer is designed as a dense-to-sparse scheme, as shown in Fig. 3.
Specifically, dense motion is first extracted with an off-the-shelf and latest motion estimation model Unimatch (unimatch), which utilizes a transformer-based multi-task framework to solve the unified dense correspondence matching problem. Without loss of generality, we denote the key frame and inter frame as I and P with the dimension of $3\times H\times W$, the motion extraction can be denoted as,
(1) $\textbf{f}_{d}=\Theta(\textbf{I},\textbf{P})$
To select the primary motion trajectories under the consideration of both its appearance and intensity, the flow edge is first extracted from the dense flow to distinguish the boundary of moving regions, then a topological-distance watershed map is created (watershed) to sample sparse flow in the center areas of moving regions, finally Non-maximum Suppression (canny) is leveraged to obtain the key-points and corresponding motion mask (cmp). Meanwhile, a uniform grid-sample is also utilized to ensure the coverage of the entire motion field. In this way, the dense motion is transformed to sparse motion and motion mask,
(2) $\textbf{f}_{s},\textbf{m}_{s}=\Phi(\textbf{f}_{d})$
where $\Phi$ denotes motion sampling process. $\textbf{f}_{s}\in\mathbb{R}^{H\times W\times 2}$ and $\textbf{m}_{s}\in\mathbb{R}^{H\times W\times 1}$ denote the sparse motion and the motion mask respectively.
Then, a sparse flow compression is performed to transform sparse motions to compact motion tokens, where the motion feature is first obtained by a flow feature extractor $\textbf{e}_{f}$,
(3) $\textbf{y}_{f}=\textbf{e}_{f}(concat[\textbf{f}_{s},\textbf{m}_{s}])$
where $\textbf{y}_{f}$ denotes the flow feature and concatconcat denotes the concatenation operation. Herein, this flow feature extractor is designed as a U-Net (unet) like structure with up-sampling, down-sampling and short-cut connections. Specifically, five down-sample layers are followed by five symmetrically designed up-sample layers. Each layer has a rescaling factor of 2 and the outputs of each down-sample layer are concatenated to the corresponding up-sampling layer. Then, two motion token vectors are derived by two vectorizers $\textbf{v}_{w}$ and $\textbf{v}_{b}$,
(4) $\textbf{w}=\textbf{v}_{w}(\textbf{y}_{f})$
(5) $\textbf{b}=\textbf{v}_{b}(\textbf{y}_{f})$
where $\textbf{w}\in\mathbb{R}^{N_{v}\times 1}$ and $\textbf{b}\in\mathbb{R}^{N_{v}\times 1}$. $N_{v}$ denotes the number of tokens in each vector. Here, the vectorizers are designed as cascaded convolutional layers and Generalized Divisive Normalization layers (factorized). Specifically, seven convolution layers with down-sample factor of 2 and one convolutional layers with down-sample factor of 3 are used to transform the spatial size of 384 to 1. In this way, high-fidelity dense motions are transformed to very compact motion tokens, which served as high-level semantic motion hints, through the guidance of intermediate sparse motions. Such a process could exploit the compact motion pattern prior in motion flows, benefiting both high compressibility of transmitted features and decent accuracy of subsequent motion recovery.
3.3. Dense Motion Recovery
Figure 4. The detailed structure of dense flow recovery module.
At the decoder side, the dense motion is recovered from the decoded motion tokens with the assistance of reconstructed key frame by a dense motion recovery module, as shown in Fig. 4. Specifically, the reconstructed key frame is denoted as $\hat{\textbf{I}}$, and it is first fed into an image feature extractor $\textbf{e}_{i}$,
@@ 3.3. Dense Motion Recovery
where $\textbf{y}_{i}\in\mathbb{R}^{N_{v}\times H\times W}$ denotes the image feature. Here, the image feature extractor is designed as a U-Net like structure, where five down-sample layers are followed by five symmetrically designed up-sample layers, and each layer has a rescaling factor of 2 and the outputs of each down-sample layer are concatenated to the corresponding up-sampling layer. The decoded motion token vector $\hat{\textbf{w}}$, $\hat{\textbf{b}}$ and the image feature are processed by a motion predictor $\textbf{g}_{f}$,
@@
where $\hat{\textbf{m}}_{f}\in\mathbb{R}^{H\times W\times 2}$ denotes the recovered dense motion flow and $\cdot$” denotes channel-wise multiplication. Here, the flow generator is designed with the same structure as the image feature $\textbf{e}_{i}$.
3.4. Flow-driven Generator
In previous GVC schemes, generators are usually crafted following the structures of Generative Adversarial Network (gan) or Variational Auto-Encoder (vae). The resulting GVC methods can only address single scenario application with one trained model, limiting their generalizability to more diverse contents. Recently, the Diffusion Model (diffusion) is proposed to perform generation with the denoising process and exhibits the outstanding capability of learning complex data distribution. It can be integrated with versatile conditions for tailored generative vision tasks (diffusion-review). Motion trajectory is also leveraged as the control signal for interactive image-to-video generation (motion-i2v; motionctrl; mofa; dragnuwa; tora). However, they are not yet explored in GVC framework. In this paper, a pre-trained SVD (svd) and motion adaptor (mofa) are utilized to ensure the robustness of motion-driven generation and its generalizability across various contents.
Following the practice in (mofa), the recovered dense motion $\hat{\textbf{m}}_{f}$ is fed into the motion-adaptor and used to warp the multi-scale features of reconstructed key frame, which are further fused with SVD encoder features. Specifically, a pre-trained SVD encoder $\textbf{E}_{ref}$ is used as reference encoder to extract the key frame features. Then, the features are warped by the recovered dense motion $\hat{\textbf{m}}_{f}$. The warped features are further fused to a trainable SVD encoder $\textbf{E}_{fus}$, which takes the noisy key frame as input and its features are finally served as denoising condition for the SVD decoder. The process can be formulated as,